A particle moves so that its position vector ˜r at time t is = ˜r ˜a coswt +˜bsinwt, where w is a constant and ˜a and ˜b are constant vectors. Show that (a) ˜ ˜ r × r˙ is independent of t, ˜


1
Expert's answer
2021-04-27T01:33:41-0400

(a)


"|\vec a|=|\vec b|=1, \vec a\perp\vec b""\vec r\cdot \vec r=(\vec a\cos\omega t+\vec b\sin \omega t)\cdot(\vec a\cos\omega t+\vec b\sin \omega t)"

"=\cos^2\omega t+\sin^2\omega t=1"


(b)


"\vec v=\dfrac{d\vec r}{dt}=-\omega\vec a\sin\omega t+\omega\vec b\cos \omega t"

"\dfrac{d\vec v}{dt}=-\omega^2\vec a\cos\omega t-\omega^2\vec b\sin \omega t"




"=-\omega^2\vec r"

The acceleration is everywhere towards the origin and proportional to "\vec r."